Circular No. 6162

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

POSSIBLE SATELLITES OF SATURN M. K. Gordon, C. D. Murray and K. Beurle, Queen Mary and Westfield

College, University of London, report on the possible existence of additional satellites of Saturn in the Voyager 2 data. They also quote S. P. Synnott as noting that the seven satellite candidates discussed by him on IAUC 3651, 3656, and 3660 should be reduced to three (S/1981 S 7, 10, and 11), together with a very faint fourth one, designated S/1981 S 14, observed on 1981 Aug. 23.9955 UT around 2.0 × 10^{5} km from Saturn. Summarized below are the 1981 Aug. UT dates of the new images (not the ‘earth received’ times), approximate planet-centered distances (in units of 10^{5} km) and longitudes (in deg), assuming circular orbits in the equatorial plane (in J2000.0 coordinates fixed on 1981 Aug. 21.2), and estimated apparent visual geocentric opposition magnitudes: S/1981 S 15, 21.2439, 1.74, 222, 18; S/1981 S 16, 21.2789, 2.20, 290, 18; S/1981 S 17, 22.1928, 2.31, 282, 18; S/1981 S 18, 23.9050, 1.85, 206, 24; S/1981 S 19, 23.9866, 1.86, 109, 22. Pointing-error corrections in the Voyager database for S/1981 S 18 and 19 could not be corrected; consequently, the error bars for these two features are on the order of two to three thousand km. The orbital radii above are particularly sensitive to the assumption that the possible satellites are on orbits in the equatorial plane. Orbits with moderate inclinations (1^{o}– 1^{o}.5) could produce changes of several thousands of kilometers in the radial positions. Based on dynamical considerations, S/1981 S 16 and 18 may be separate observations of a single small satellite coorbital with Mimas (Saturn I) with an orbital inclination comparable to that of Mimas.

AL COMAE BERENICES W. Pych and A. Olech, Warsaw University Observatory, write: “Time-

series observations of AL Com were collected with the 0.60-m telescope (+ Tek512 CCD camera) on three nights (Apr. 10.026–10.104, 10.805–10.896, and 11.783–11.994 UT) at the Ostrowik Observatory of Warsaw University. Observations were made in the Cousins I (first night) and V bands. We confirm the presence of coherent periodic light variations (cf. IAUC 6157). The best period for our entire data set is 81.5 ± 0.2 min. Also, the first harmonic period (40.8 min) is clearly detectable. The peak-to-peak amplitude in the V and I bands was ∼ 0.05 and 0.06 mag, respectively.”

1995 April 14 Daniel W. E. Green